OpenStudy (anonymous):

Solve for x: sin^(-1)(1-x)-2sin^(-1)x= π/2

OpenStudy (anonymous):

sin^(-1) is inverse sine, right?

OpenStudy (anonymous):

\[\sin^{-1}(1-x)-2\sin^{-1}x= π/2\]

OpenStudy (anonymous):

It is sin inverse

OpenStudy (anonymous):

Wolframalpha gives the answer as x=0

OpenStudy (anonymous):

Steps please?

OpenStudy (anonymous):

take sin inverse (1-x) as alpha and sin inverse x as theta sin(alpha-2theta)=sin(pi/2) sinalphacos2thta-cosalphasin2theta=sinpi/2 apply the corresponding values then u vl get the answr

OpenStudy (anonymous):

ok i will try ThanQ ;) here comes your medal